OpenStudy (anonymous):

What is the denominator of the simplifed form of 1/2 - 3/4x-2? a. 2 b. 4 c. x-2 d. 2x-1

OpenStudy (anonymous):

\[\frac{1}{2} - \frac{3}{2(2x - 1)} = \frac{2x -1 -3}{2(2x - 1)} = \frac{2(x-4)}{2(2x - 1)}\] So what will be left after cancelling two in the denominator???